OpenStudy (anonymous):

What is the reference angle for sec -270 degrees?

OpenStudy (jdoe0001):

keep in mind that cos(-a) = cos(a) 1/cos(-270) = 1/cos(270) that's \(\bf \cfrac{3\pi}{2}\)